OpenStudy (anonymous):

Find the derivative of f at x. That is, find f '(x). f(x) = 7x + 8; x = 5

OpenStudy (therealmeeeee):

f(x) = x^6 • 7e^(7x) + e^(7x) • 6x^5 = e^(7x) (7x^6 + 6x^5) hope this help